½ of 4 whole 2/7 is                                       
½ of 4 whole 2/7 is                                       
½ of 4 whole 2/7 is                                       

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1⁄2 of 4 whole 2/7 is 15/7.

First we have to convert the mixed fraction into improper fraction 4 whole 2/7= 30/7 Then, 1⁄2 × (30/7)

= 15/7

Thanks